Find the cos 0 (theta) if csc 0(theta) = 11 square root of 2 divided by 12. (I don't have the symbols on my computer to write this exactly like on the page, but it is 11 by square root of two, over 12).
Answer by stanbon(57307)

Find the cos (theta) if csc theta) = 11 square root of 2 divided by 12
------------------
General Form: csc(theta) = r/y
-------
Your Problem:
csc(theta) = [11sqrt(2)]/12 , so r = 11sqrt(2) and y = 12
Find x: x = sqrt(r^2-y^2) = sqrt[121*2 + 144] = sqrt(386)
------------------
Since cos(theta) = x/r, cos(theta) = sqrt(386)/11sqrt(2) = sqrt(193)/11
========================================
Cheers,
Stan H.